What is Average Age of Accounts Payable for the period ending March 31 2016 based on the data provided in the following Tables?

A) 58

B) 90

C) 99

D) 47
Answer #1

Average age of account payable= average account payable× 365/ (cost of constructions)

Cost of construction is Total direct materials cost (cost of Goods sold)

Average account payable= beginning account payable + ending account payable/2

Average account payable= (560,000+475,000+ 1,270,000+ 400,000) /2= 1, 352,500

Average age of account payable= 1,352,500× 365/9,000,000 = 58